OpenStudy (anonymous):

Simplify the expression (5x²+2x+4/x⁴-16) / (25x²+20x+4/x² -2x)

OpenStudy (anonymous):

\[\frac{ 5x²+2x+4}{ x^4-16 } \div \frac{ 25x^2+20x+4 }{ x^2-2x }\] do u know how to factor this

OpenStudy (anonymous):

I kind of remember the idea but I forgot the proper way to do it :/

OpenStudy (anonymous):

so lets see a simple one take \[x ^{4}-16\] = \[(x ^{2}-4)(x ^{2}+4)\]

OpenStudy (compassionate):

Still need help?

OpenStudy (luigi0210):

When you are diving, you flip the function on the right and switch it to multiplying.

OpenStudy (luigi0210):

No ansurs here, sorry

OpenStudy (anonymous):

\[(\frac{ 5x²+2x+4 }{ x⁴-16 }) (\frac{ x ^{2}-2 }{ 25x²+20x+4 })\] Like this? And do I factor these now? For x²-2 I got (x+2)(x-1) ?? Is this correct? Or do I have to cross multiply first and combine like terms? I know this must me alot simpler than Iam making it, It's just been a while since I took a math course.

OpenStudy (luigi0210):

I thought it was x^2-2x?

OpenStudy (anonymous):

You're correct. Sorry, it's been a long day -_-

OpenStudy (luigi0210):

And that would make it simpler wouldn't you agree?

OpenStudy (anonymous):

x(x-2) ??

OpenStudy (luigi0210):

yup!

OpenStudy (anonymous):

So how do I factor 25x²+20x+4? Or do I just keep it like this?

OpenStudy (luigi0210):

Factor it, it might help.

OpenStudy (anonymous):

As I implied. I'm not so sure how I'd factor it This part is where I'm stuck and my book has not come in yet so I don't even have a reference. :/ How would I start on factoring it? :O

OpenStudy (luigi0210):

It factors to (5x+2)(5x+2)